# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

### v1

@@ C.3 Generation Settings
SamplingParams(    temperature=0,    repetition_penalty=1.1,    max_tokens=128,    top_p=1.0,    top_k=-1,    stop_token_ids=[eos_token_id] if eos_token_id is not None else [],    stop=[
@@
Appendix D Cross-lingual Alignment
@@
Appendix E Computational Budget
All experiments were conducted on a single NVIDIA H100 GPU, except for the full fine-tuning setup, which required two H100 GPUs due to its higher memory and compute demands. Each model was trained for approximately 100 million tokens, corresponding to about 6-8 GPU hours per run for parameter-efficient methods and roughly 10 GPU hours for full fine-tuning.
Appendix F Licenses and Attribution
All experiments were conducted using publicly available models and datasets released under permissive or research-oriented licenses. The Llama-3.1-8B model is distributed under the Meta Llama 3 Community License Agreement, which permits research use and redistribution with attribution. Mistral-Nemo-12B is released under the Apache License 2.0, allowing open use and modification. Language-specific token data were obtained from the GlotCC corpus (CC0 for metadata, Common Crawl terms for text content). Evaluation datasets including FLORES, BELEBELE, SIB200, and OPUS were used in accordance with their respective research licenses, and English benchmarks (ARC, Hellaswag, PIQA, Winogrande, MMLU) were accessed under standard academic usage terms.

### v2

@@ E.2 Generation Settings
SamplingParams(
    temperature=0,
    repetition_penalty=1.1,
    max_tokens=128,
    top_p=1.0,
    top_k=-1,
    stop_token_ids=[eos_token_id] if eos_token_id is not None
            else [],
    stop=[
@@
Appendix F FFN Component Ablations
To understand which parts of the FFN contribute most to language-specific adaptation, we perform ablations by fine-tuning individual or combinations of the up-, gate-, and down-projection matrices of the identified language subnetworks. Table 39 reports the resulting performance across a subset of six languages.
FFN component mt af sk mk lv ne Avg.
Full subnetwork size 37.0M 31.7M 23.8M 21.4M 39.5M 56.8M
base 30.04 37.80 33.43 33.32 29.88 27.06 31.92
ffn (up+gate+down) 39.01 42.99 35.56 34.98 32.20 32.65 36.23
up 37.17 39.98 33.48 35.27 30.83 30.69 34.57
gate+up 38.14 42.18 34.81 34.13 31.28 27.23 34.63
down 34.99 40.60 35.31 36.57 30.36 28.66 34.42
up+down 37.13 44.11 35.00 35.28 33.71 29.72 35.83
Table 39: Average performance of fine-tuning different FFN components of Llama-3.1-8B across six languages. up, gate, and down denote the identified language-specific components within the FFN up-, gate-, and down-projection weight matrices, respectively. Fine-tuning a single matrix corresponds to one third of the full language subnetwork size, while fine-tuning two matrices corresponds to two thirds.
Appendix G Correlation with Language Resource Availability
To examine whether the observed improvements from targeted fine-tuning are systematically related to language resource availability, we analyze correlations using two complementary metrics.
CC-100 Corpus Size.
We use the CC-100 corpus size (in MB) as a proxy for the relative digital resource availability across languages (Conneau et al., 2020). While these values represent web-crawled text from a specific snapshot, they serve as reasonable approximations of the relative proportions of available internet data per language. The ratios between languages remain informative even if absolute quantities have changed over time.
Joshi Resource Class.
We additionally employ the taxonomy proposed by Joshi et al. (2020), which classifies languages into six categories (0–5) based on both labeled and unlabeled data availability. This discrete classification considers multiple factors beyond raw corpus size, including the availability of annotated datasets and representation in NLP research. For our target languages, resource classes range from 1 (Maltese) to 3 (Estonian, Lithuanian, Latvian, Slovak, Slovenian).
LLaMA-3.1-8B Aya-Expanse-8B Mistral-NeMo-12B
Figure 14: Correlation between monolingual improvement scores and language resource metrics. Top row: CC-100 corpus size in MB, used as a proxy for relative internet data availability (Conneau et al., 2020). Bottom row: Resource class from the Joshi et al. taxonomy (Joshi et al., 2020). Pearson correlation coefficients are shown; * indicates $p<0.05$. Performance gains from targeted fine-tuning are highest for lower-resource languages and diminish as resource availability increases, suggesting limited marginal benefit for languages already well represented during pre-training.
Appendix H Cross-lingual Alignment
We additionally provide detailedper-language cosine similarities between parallel sentences before and after fine-tuning for all models. These results complement Section 5 by illustrating how targeted subnetwork adaptation affects cross-lingual alignment across languages.